Prev Next

SW&IT Meeting Minutes 2020-04-13

Attendees: Carl, David, Igor, LeRoy, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200406
LR20200406 - Bec, Matthieu 40.0
LR20200406 - Holmberg, Carl 40.0
LR20200406 - Edmeades, David 40.0
LR20200406 - Sola, Igor 40.0
LR20200406 - Hooper, Stephen 80.0
LR20200406 - Zhang, Yang 40.0
LR20200406 - Zhang, Xianyu 51.0

News from management

  • IT reform discussions (SB)

Previous Action Items

Owner Action Due Date Completion DateSorted ascending
IS/DE create a new Luci simulator for ROWG purpose (+ work out resources available on the VM)    
DE/XZ Xianyu's IDL personal license    
MB SOUL/UAO follow-up    
IS dig-out how LUCI computes its WCS    

Open topics

  • Luci simulator for ROWG/UI tests with new UHD screens
  • AO data system (document from XZ)

Team reports

MB

  • Last week
    • OC, Management
    • ROWG#5
    • DMS AdSec act #
    • Activity sheet for CV
    • SOUL / UAO conversation with Arcetri - they don't have the resources. Issue LN/Argos. We don't have the LN s/w, we could potentially fix Argos. In Julian's hand to iterate.
  • This week
    • cobbler / kickstart
    • ims f30
    • student ad
    • DMS weather loss, websockets, AdSec, etc.
    • Jira ?
    • Shark-VIS follow-up
    • LBC remote access
    • DMS APi demo for Josh
    • indi-core check out with CH

DE

  • Pretty well convinced we won't be affected by Friday's power work
  • Learned that vibration laptop has been retasked; working with Peter on new spec
    • With the scope creep we've experienced, this is good since the $200 one I chose before was not up to what people were asking of it
  • Laura monitor/dock issues mostly solved but watching
  • Julian ISP "advanced security" option blocks access to VPN.
  • New persistent Zoom role account for AO group
  • Software installs for DR. Will follow up this week
  • Avtech server room monitoring hardware/software looks promising; shall we order a test bundle?
  • Tested some beginner security on Zoom rooms
    • Discovered that not everyone was logging into their UA Zoom account
    • Should we amplify this to all of LBTO, in case we do need to quickly start adding some restrictions?
  • Mac Mini finally arrived!
Next
  • Work on pointing Veeam to new VM Host
  • Open ticket with Raritan to see if we can figure out performance problems of KVM
    • Stephen, can we talk about the issues you've encountered so I can speak intelligently to them?
  • Luci.luci upgrade specifics/are there any new servers that have PCI risers as options?
  • Laptop specs
    • PG/vibration
    • Laura

SH

  • vm1.tucson upgraded (new wiring, re-rack, upgrade VM)
  • ovirt (media not working on R410s)
  • Tucson VPN routes (need to re-wire Tucson VPN -- maybe use spare VPN)
  • FTP for MGIO cameras. Talk to Leroy about changing camera location.
  • ROWG and other meetings
  • AO telemetry
  • User issues (sound, network, VPN)

next:

  • ovirt (again)
  • AllSky (again)
  • Tucson CR work (computer install for ROWG)
  • SSL for LBTC
  • ROWG purchases

XZ

  • Last week
    • PB SW document and meeting
      • requirements to SOUL SW for PB SW
      • PPT presentation for SOUL/UAO SW
      • pre-Arcetri meeting
      • discuss with Arcetri
    • AO telemetry data
    • write and send out AO telemetry document
    • Actuators status analysis
      • setup environment and run test
      • current and postion disable detection
    • Meetings
  • This week
    • PB action items
    • PB SW/HW
    • Actuators status analysis

YZ

  • Last week:
    • IT#7854, the LBC Web GUI sometimes loses track of the true internal state of LBC. Examined the LBC code more closely, have some hypotheses, putting it off until LBC log files are accessible.
    • IT#8021, LBC rotator trailing issues. Looked at some of the log file entries. Seems to be fine. So far no real clues (other than some timing issues discovered by Olga, but unknown if they matter).
    • MODs Python program/Pyraf study. Pyraf was used to communicate the DS9 on screen coordinates with MODS Python program, so there may be a solution to remove Pyraf dependency.
    • Free Mathematica (unlimited time) on the Raspberry Pi (the official Raspbian OS). Configured my old Raspberry Pi 3 and was able to load Andrew's Mathematica notebook. But with lots of questions unanswered in his notebook.
    • Setup UA zoom account.
  • This week:
    • Continue IT#8021, LBC rotator trailing issues. Current plan is to study an old issue IT#3346, which contains lots of information regarding LBC rotator trajectory related issues.
    • Study DS9 XPA details to see if we can directly communicate with MODS Python program (Matthieu provided some references).
    • Will examine the TMS Mathematica notebook more closely and draft some questions.
    • AO-Cravat work on Python/C Strehl code reorganization.
    • Discuss with Olga about the LBC/MODS fits header information.
    • Safety training.

CH

  • Last week.
    • ARGOS, initial work to isolate GUI from back end.
    • AHS, finish repackaging INDI library and commit. Still issues with Python module install.
  • This week.
    • ARGOS, working on fixing Jenkins build, create custom build script for GUI front end.

IS

Last week:
  • PIT
    • Removed Darkest option from Sky Background menu and PDF and made conversion of older releases for backward compatibility sake.
    • Revisit declination logic and added an error message for a limit case that wasn't handled in the error messages although it was in the visibility display.
    • Made new release INAF PIT 2.1003.0 test release
  • OT
    • Worked with Andrew on OT bugs and clarifying testing issues.
  • LUCI
    • Catch up with detector channel issue emails
    • Refreshing some of the AFC development done with Dave T.
    • CalUnit Web I/O got to my home. Will look into configuration probably at the end of week or next week.

This week:
  • PIT
    • If everything goes well do final release for INAF
    • Keep working on the PIT admin tool backend.
  • OT
    • work with Andrew on current test OT release issues
  • LUCI
    • install latest GEIRS in simulation environment
    • maybe install the current production database, might need more space, already touch base with David E. on this.
* Look into LUCI WebIO device configuration. * Look into LUCI WebIO device configuration.
Topic revision: r16 - 13 Apr 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback